# Neutrino Oscillations in Core-Collapse Supernovae and Neutron Star Mergers

### v1

@@ 4.4 Matter effects
Clearly there is some validity to the rotating-frame argument, and it is certainly a boon not to have $\lambda$ determining the time step in numerical calculations. But in light of the many advances of recent years, a comprehensive reassessment of matter effects is now warranted. A thorough analysis will need to account for the matter-flux potential as well. Although this term has almost always been neglected, it is not obviously insignificant [130]. The structure of the term puts it on par with the anisotropic part of the neutrino–neutrino potential: compare the factors of $\bm{\hat{q}}\cdot\bm{v}_{m}$ and $\bm{\hat{q}}\cdot\bm{D}_{1}$ in Eq. 4. The matter-flux potential can be competitive with the neutrino–neutrino potential even though $|\bm{v}_{m}|\ll 1$.
@@ 4.5 Turbulence and equilibrium
In an earlier era of research on collective oscillations, it was hypothesized that neutrino–neutrino forward scattering might produce neutrino distributions of such fine structure—“developing modes of higher and higher multipolarity”—that flavor equipartition would effectively be achieved [131]. Years later, this suggestion seems rather prescient. It contains two ideas that are now considered essential. An abundance of numerical evidence shows that flavor instabilities, at least SFI and FFI, generally coincide with a cascade in phase space, whereby neutrino flavor distributions form features on progressively smaller scales [95, 132, 133, 134, 135, 136]. This phenomenon is a form of turbulence facilitated by the nonlinearity of neutrino refraction in dense environments. A fluctuating or quasi-steady asymptotic state is reached, which is sometimes referred to as flavor equilibrium (though it not a state of equipartition except under special circumstances).
@@ 5.2 Quantum closures
More recent work has sought to formulate suitable quantum closures for moment-based quantum kinetics. Initial studies using ad hoc closure prescriptions found enough success in replicating the results of linear stability analysis and nonlinear evolution to warrant further investigation [171, 172, 173, 174]. The challenges now are to construct generally applicable and theoretically sound quantum closures [175, 176] and to ensure full compatibility between quantum moments and the computational methods of astrophysical simulations [177].
@@ 5.3 Coarse-grained theories
An alternative asymptotic-state method replaces $[H_{\bm{q}},\rho_{\bm{q}}]$ in the QKE with the quantum BGK approximation $\Gamma^{\textrm{a}}(\rho^{\textrm{a}}_{\bm{q}}-\rho_{\bm{q}})$, where $\Gamma^{\textrm{a}}$ is the relaxation rate and $\rho^{\textrm{a}}_{\bm{q}}$ is the asymptotic state [161]. This quantum relaxation-time approximation can be compared with the collisional relaxation-time approximation in Sec. 3.2. As in the two-step procedure, oscillations are assumed to bring neutrinos toward local asymptotic states. The intent is to allow for finite rates of relaxation toward these states, rather than imposing instantaneous relaxation. Qualitative agreement has been found when comparing BGK and quantum kinetic solutions in a model with neutrino injection from the inner boundary of a radial shell [161].
Regarding the calculation of asymptotic states, the optimal choice would be to run the actual isolated-box QKE calculation for every unstable region. Because this is computationally prohibitive, approximate options are being explored (Sec. 4.5). Among these are various empirical formulas motivated by QKE calculations [142, 141, 143, 144], including the two shown in the right panel of Fig. 5, which were adopted in Ref. [165]. Other methods use machine learning to develop the mapping from unstable flavor configurations to asymptotic states [145, 146, 143]. Irrespective of which option is selected, boundary conditions must be chosen whenever a small region is plucked out and isolated from the rest of the astrophysical environment. Scale separation of the kind illustrated in Fig. 6 motivates periodic boundary conditions, as adopted in the references above. Asymptotic states differ when Dirichlet boundary conditions are used instead [147, 148].
@@ ACKNOWLEDGMENTS
We acknowledge Sajad Abbar, Ryuichiro Akaho, Jakob Ehring, Julien Froustey, and Zewei Xiong for providing figures and simulation data. LJ is supported by a Feynman Fellowship through LANL LDRD project number 20230788PRD1. SR is supported by the National Science Foundation under Grant No. 2412683. MRW acknowledges support of the National Science and Technology Council, Taiwan under Grant No. 111-2628-M-001-003-MY4, the Academia Sinica under Project No. AS-IV-114-M04, and the Physics Division of the National Center for Theoretical Sciences, Taiwan.

### v2

@@ 4.4 Matter effects
Clearly there is some validity to the rotating-frame argument, and it is certainly a boon not to have $\lambda$ determining the time step in numerical calculations. But in light of the many advances of recent years, a comprehensive reassessment of matter effects is now warranted [142, 31, 42, 143]. A thorough analysis will need to account for the matter-flux potential as well. Although this term has almost always been neglected, it is not obviously insignificant [144]. The structure of the term puts it on par with the anisotropic part of the neutrino–neutrino potential: compare the factors of $\bm{\hat{q}}\cdot\bm{v}_{m}$ and $\bm{\hat{q}}\cdot\bm{D}_{1}$ in Eq. 4. The matter-flux potential can be competitive with the neutrino–neutrino potential even though $|\bm{v}_{m}|\ll 1$.
@@ 4.5 Turbulence and equilibrium
In an earlier era of research on collective oscillations, it was hypothesized that neutrino–neutrino forward scattering might produce neutrino distributions of such fine structure—“developing modes of higher and higher multipolarity”—that flavor equipartition would effectively be achieved [145]. Years later, this suggestion seems rather prescient. It contains two ideas that are now considered essential. An abundance of numerical evidence shows that flavor instabilities, at least SFI and FFI, generally coincide with a cascade in phase space, whereby neutrino flavor distributions form features on progressively smaller scales [107, 146, 147, 148, 149, 150]. This phenomenon is a form of turbulence facilitated by the nonlinearity of neutrino refraction in dense environments. A fluctuating or quasi-steady asymptotic state is reached, which is sometimes referred to as flavor equilibrium (though it is not a state of equipartition, with an equal number of neutrinos in each flavor, except under special circumstances).
@@ 5.2 Quantum closures
More recent work has sought to formulate suitable quantum closures (i.e., angular-moment closures that include information about flavor coherence) for moment-based quantum kinetics. Initial studies using ad hoc closure prescriptions found enough success in replicating the results of linear stability analysis and nonlinear evolution to warrant further investigation [185, 186, 187, 188]. The challenges now are to construct generally applicable and theoretically sound quantum closures [189, 190] and to ensure full compatibility between quantum moments and the computational methods of astrophysical simulations [191].
@@ 5.3 Coarse-grained theories
An alternative asymptotic-state method replaces $H_{\bm{q}},\rho_{\bm{q}}]$ in the QKE with the quantum BGK approximation $\Gamma^{\textrm{a}}(\rho^{\textrm{a}}_{\bm{q}}-\rho_{\bm{q}})$, where $\Gamma^{\textrm{a}}$ is the relaxation rate and $\rho^{\textrm{a}}_{\bm{q}}$ is the asymptotic state [175]. This quantum relaxation-time approximation can be compared with the collisional relaxation-time approximation in Sec. 3.2. As in the two-step procedure, oscillations are assumed to bring neutrinos toward local asymptotic states. The intent is to allow for finite rates of relaxation toward these states, rather than imposing instantaneous relaxation. Qualitative agreement has been found when comparing BGK and quantum kinetic solutions in a model with neutrino injection from the inner boundary of a radial shell [175].
Regarding the calculation of asymptotic states, the optimal choice would be to run the actual isolated-box QKE calculation for every unstable region. Because this is computationally prohibitive, approximate options are being explored (Sec. 4.5). Among these are various empirical formulas motivated by QKE calculations [156, 155, 157, 158], including the two shown in the right panel of Fig. 5, which were adopted in Ref. [179]. Other methods use machine learning to develop the mapping from unstable flavor configurations to asymptotic states [159, 160, 157]. Irrespective of which option is selected, boundary conditions must be chosen whenever a small region is plucked out and isolated from the rest of the astrophysical environment. Scale separation of the kind illustrated in Fig. 6 motivates periodic boundary conditions, as adopted in the references above. Asymptotic states differ when Dirichlet boundary conditions are used instead [194, 161, 162].
@@ ACKNOWLEDGMENTS
We acknowledge Sajad Abbar, Ryuichiro Akaho, Jakob Ehring, Julien Froustey, and Zewei Xiong for providing figures and simulation data. We are grateful to Basudeb Dasgupta, Damiano Fiorillo, Gabriel Martínez-Pinedo, Georg Raffelt, Günter Sigl, Irene Tamborra, and Zewei Xiong for comments on the manuscript. LJ is supported by a Feynman Fellowship through LANL LDRD project number 20230788PRD1. SR is supported by the National Science Foundation under Grant No. 2412683. MRW acknowledges support of the National Science and Technology Council, Taiwan under Grant No. 111-2628-M-001-003-MY4, the Academia Sinica under Project No. AS-IV-114-M04, and the Physics Division of the National Center for Theoretical Sciences, Taiwan.